# Traditional Methods Outperform Generative LLMs at Forecasting Credit Ratings

### v1

@@ Abstract.
Large Language Models (LLMs) have been shown to perform well for many downstream tasks. Transfer learning can enable LLMs to acquire skills that were not targeted during pre-training. In financial contexts, LLMs can sometimes beat well-established benchmarks. This paper investigates how well LLMs perform in the task of forecasting corporate credit ratings. We show that while LLMs are very good at encoding textual information, traditional methods are still very competitive when it comes to encoding numeric and multimodal data. For our task, current LLMs perform worse than a more traditional XGBoost architecture that combines fundamental and macroeconomic data with high-density text-based embedding features. The code from this paper is provided [^1] https://github.com/FelixDrinkall/credit-ratings-project.
Forecasting, NLP, LLM, Credit Ratings, Generative, XGBoost, Clustering, Multimodal
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; pre-print; July 10th, 2024
1. Introduction
@@
Corporate credit ratings indicate a borrower’s ability to service its debt obligations and are a forward-looking measure of a company’s health (Baresa et al., 2012). A company’s credit rating is significant since it affects the cost of raising capital, which in turn could finance future infrastructure to increase revenue or profitability. An optimistic rating can result in a virtuous cycle whereby it is easier to raise money and grow the business (Cho et al., 2020), and a pessimistic rating can result in a vicious cycle in which competition can grow faster due to cheaper debt obligations. Knowing which cycle a company may enter can be advantageous to investors. Many major funds are also not allowed to own sub-prime assets, which makes forecasting a drop in credit rating very important so that the fund has more time to divest from the asset, which could result in a higher close price.
@@
Figure 1. Example of the best-performing feature-type - high-density clustering (Drinkall et al., 2022). Each dot represents a sentence, and the colored dots represent sentences that exist in high-density regions of embedding space.
@@
Recently, there has been a surge of interest in text-based forecasting (Xu and Cohen, 2018; Yang et al., 2020; Nie et al., 2024). One reason for this trend is the progress that has been made in text modelling in general (Wei et al., 2022; Touvron et al., 2023). Given that financial news is often first disseminated through written or spoken communications (Boulland et al., 2016), rather than in numeric or tabular formats, there has been a hope that important information can be included in models sooner than was possible without using linguistic information. Another reason is that language can provide relevant context and forward-looking information, whereas financial numeric reporting alone is inherently retrospective. Contained within a company’s filings, text can provide insights about the future strategic direction of the company as well as historical information.
@@
The majority of text-based forecasting research has been directed at analyzing short-sequence textual content, drawing primarily from sources such as social media (Xu and Cohen, 2018), news articles (Zhang et al., 2018), and analyst recommendations (Rekabsaz et al., 2017). In contrast, many fundamental financial documents, such as company filings, earnings call transcripts, and patents, are very long. Considering that the shorter texts often serve as summaries, reflections or commentaries on the detailed primary sources and that speed of information acquisition is essential in finance (Rzayev et al., 2023), there must be more focus on text-based forecasting for longer text sequences. This paper evaluates the most effective ways to model longer text sequences within a text-based forecasting task.
@@
Linked to the recent progress of LLMs, there has been growing interest in applying language models in a wide variety of downstream applications (Kaddour et al., 2023). It has been shown that as generative LLMs scale, they acquire abilities that were not present in smaller LLM variants (Wei et al., 2022), such as modular arithmetic (Srivastava et al., 2023), NLU (Hendrycks et al., 2021), commonsense reasoning (Lin et al., 2022), fact-checking (Rae et al., 2022) and so on. These abilities have been impressive, but it is best not to be over-optimistic. The lack of training data transparency associated with some of the best-performing language models means that we cannot be certain whether some of the performance gains are due to the memorisation of benchmarks being in the training datasets (Bender et al., 2021; Sainz et al., 2023; Xu et al., 2024; Balloccu et al., 2024). Generative language models also seem to have a mediocre understanding of concepts like negation and complex logical reasoning (Kassner and Schütze, 2020; Lorge and Pierrehumbert, 2023; Huang and Chang, 2023; Truong et al., 2023). These limitations in the capabilities of language models could prove to be very consequential in financial contexts. In this paper, we test generative language models on a complex linguistic task, which has never been fully solved by human experts: credit rating forecasting. We show that while language models encode text-based information very well, they are not good at incorporating numeric information, and underperform a boosting-tree baseline.
1.1. Contributions
@@
  • A benchmark of different techniques for encoding long-sequence text in a forecasting task.
2. Related Work
2.1. Text-based forecasting
2.1.1. Encoding Text for Forecasting
@@
The predominant approach in text-based forecasting has focused on the extraction of interpretable features like sentiment and uncertainty scores (Song and Shin, 2019; An et al., 2023). Rule-based sentiment using diverse lexicons has dominated the literature (Mohammad, 2020; Kalamara et al., 2022; Barbaglia et al., 2023). Lexicons tailored to specific domains generally surpass broader lexicons in predictive tasks (Loughran and McDonald, 2011; Li et al., 2014). Nevertheless, lexicons overlook contextual nuance and inadequately address common linguistic phenomena like negation. To mitigate these limitations, efforts have been made to integrate more sophisticated sentiment classifiers (An et al., 2023; Ayyappa et al., 2023). However, sentiment presupposes that important information can be encapsulated within a single dimension. To avoid an overly simple and prescriptive feature set, unsupervised methods have been used in feature exploration: TF-IDF (Jones, 1972), Latent Dirichlet Allocation (LDA) (Wang et al., 2017; Kanungsukkasem and Leelanupab, 2019). However, the arrival of contemporary topic models has gradually eclipsed LDA, fostering the adoption of transformer-derived topic models into forecasting tasks (Drinkall et al., 2022).
@@
Recently, some studies have used the representations from encoder-based LLMs as features for text-based forecasting. LLMs exploit high-dimensional embeddings to capture the linguistic meaning of words (Devlin et al., 2019; Radford et al., 2019) and sentences (Reimers and Gurevych, 2019), with representation dimensionality ranging from 384 (Wang et al., 2020) to 5192 (Touvron et al., 2023). Such dimensionality poses challenges when used with smaller datasets. Nonetheless, there has been some success in incorporating these methods into text-based forecasting (Sawhney et al., 2020; Lee et al., 2023). However, the effectiveness of LLMs is often hampered by their limited context windows. Recent advancements have seen an increase in context window sizes, thanks in part to better GPU infrastructure making it computationally feasible, the implementation of attention sparsification techniques (Tay et al., 2022), and positional encoding hacks (Chen et al., 2023). There are also methods that combine the use of transformer-based LLMs with feature-based methods, such as topic clusters (Grootendorst, 2022; Drinkall et al., 2022), or emotions (Liapis and Kotsiantis, 2023).
2.1.2. Generative Multimodal Forecasting
@@
In addition to the adoption of encoder-based LLMs like BERT (Devlin et al., 2019), generative LLMs have been used in text-based forecasting tasks. Generative LLMs use masked-self attention to model text in an autoregressive manner. The GPT (Radford et al., 2019) and Llama (Touvron et al., 2023) model families are part of the generative model class. LLMs have been used as a backbone model for generative time-series forecasting models (Cao et al., 2023; Chang et al., 2023; Zhou et al., 2023; Liu et al., 2024), showing that an adapted generative language model can forecast weather, electricity and several other domains without relying on traditional text inputs. (Liu et al., 2024) used eight text-based frames in order to create a general time-series modal that could be applied to several domains, and in so doing encoded both text and numerical information in a GPT-2-small model to generate the predicted time-series.
@@
Beyond the use of text-based frames in generative forecasting tasks, GPT4MTS (Lee et al., 2023) encoded both news and time-series information before passing the concatenated input sequence through a pre-trained GPT-2 model. FinMA (Xie et al., 2023) and PromptCast (Xue and Salim, 2023) evaluated the performance of language models on stock movement prediction by converting the time-series information into natural language and prompting the language model for the predicted direction. (Yu et al., 2023) takes this further by passing exclusively text information into the prompt for a financial forecasting task. There has been little comparison between these generative methods and the more traditional discriminative methods when applied to multimodal information.
2.2. Credit Rating Prediction
@@
Research in Credit Rating Prediction (CRP) has tended to focus on predicting the absolute credit rating at time $t=0$ given the feature set $F_{t=0}$ (Li et al., 2023; Galil et al., 2023; Tavakoli et al., 2023). This approach takes the perspective of the rating agencies and is useful for identifying anomalies where the existing credit rating classification appears to be implausible or inconsistent with current financial indicators (Lokanan et al., 2019). However, predicting the absolute rating level is more simple and not as useful as predicting a future change. There is some limited research on Credit Rating Forecasting (CRF), where the target is the movement direction of the credit rating at time $t=1$. This task takes the perspective of the investor seeking to predict whether an asset is likely to be classified as more or less risky in the next time period, and is the task outlined in this paper.
@@
There have been some attempts to incorporate linguistic information into both corporate risk (Fei et al., 2015; Cao et al., 2024) and default prediction (Mai et al., 2019; Stevenson et al., 2021). Some papers have shown how text can help improve consumer credit lending (Hurley and Adebayo, 2016; Babaei and Giudici, 2024). There has also been some attempts to include textual data in CRP and CRF tasks (Chen and Chen, 2022; Muñoz-Izquierdo et al., 2022; Tavakoli et al., 2023). The majority of the existing literature uses lexicons, keywords or sentiment to encode the text (Kogan et al., 2009; Fei et al., 2015; Mai et al., 2019; Muñoz-Izquierdo et al., 2022; Chen and Chen, 2022). There have been some studies that have utilised encoder-based LLM representations (Stevenson et al., 2021; Tavakoli et al., 2023; Cao et al., 2024). There has been some work exploring how well generative models perform at assessing credit lending applications (Babaei and Giudici, 2024), and value at risk in general (Cao et al., 2024), but there has been no work benchmarking how well modern generative language models perform on a CRF task. Understanding generative LLMs’ relative strengths relative to more traditional methods is an important contribution to the existing literature.
3. Dataset
@@
In part due to the lack of large open-source or readily available datasets with temporal metadata, most of the financial text-based forecasting studies have either focused on expensive proprietary datasets, or datasets spanning 2-3 years (Xu and Cohen, 2018; Soun et al., 2022), making results hard to replicate and potentially biased to a specific time. While temporal bias in language-based tasks is hard to avoid due to limited historical data (Drinkall et al., 2024), we aim to reduce this by using a dataset spanning 23 years, enhancing the generalizability of our findings across different economic contexts. The cost and lack of transparency of large datasets have hindered progress in the field and made it harder to build on promising work due to the difficulty of replicating results.
3.1. Data Sources
All data used is either open source or available with a WRDS subscription to enable effective dataset reconstruction. The data used in this paper is derived from exclusively US-based companies.
3.1.1. Credit ratings (CR)
@@
For the credit ratings we used the Compustat Capital IQ dataset [^2] Credit Ratings: https://tinyurl.com/r4urtkc5, using Standard & Poors’ (S&P) ratings. These ratings cover the period from 1978 to 2017. S&P routinely assesses and assigns credit ratings to companies. Our paper predicts changes to the long-term credit ratings. Notably, we incorporate historical ratings from preceding quarters into our prediction models, acknowledging the distinct implications of a top-rated company (AAA) being downgraded compared to a lower-rated one (CC) experiencing a similar decline.
3.1.2. SEC filings
@@
This paper utilizes 10-Q and 10-K filings available in the SEC’s EDGAR database [^3] Filings Database: https://www.sec.gov/edgar/searchedgar/companysearch to provide both textual context. They were chosen for their consistent structure which aids homogenous feature extraction. While most of the content in these filings is comprised of indexing, tables, and introductory text, we’re interested in the parts that offer insights into a company’s future financial health. As such, we’ve focused on the Management’s Discussion and Analysis of Financial Condition and Results of Operations (MDA) section. This paper extracts the MDA sections from all SEC filings - using the SEC-API [^4] Filings API: https://sec-api.io/ - for which we had credit rating data, spanning from Q1 of 1994 to Q2 of 2017. The API itself returns cleaned text, but we clean the text further by removing the remaining HTML, links and excessive spaces.
3.1.3. Fundamental data (F)
@@
S&P emphasize two primary components in their credit rating methodology: the financial and business risk profiles (Gillmor, 2015). Whilst the text from the MDA section can provide some insight into the qualitative business risk profile, numerical fundamental data is important to assess the financial health of a company. For the fundamental data, we use the Compustat Quarterly Fundamentals dataset [^5] Fundamental Data:https://tinyurl.com/4ca8ddst. The variables selected are outlined in Appendix D. Importantly, these variables were consistently reported for all the companies under consideration. Ideally, we would incorporate a broader range of fundamental variables, but expanding the variable set would result in fewer samples with complete data, thus limiting the scope of our analysis.
3.1.4. Macroeconomic data (M)

### v2

# Forecasting Credit Ratings: A Case Study where Traditional Methods Outperform Generative LLMs

@@ Abstract
Large Language Models (LLMs) have been shown to perform well for many downstream tasks. Transfer learning can enable LLMs to acquire skills that were not targeted during pre-training. In financial contexts, LLMs can sometimes beat well-established benchmarks. This paper investigates how well LLMs perform at forecasting corporate credit ratings. We show that while LLMs are very good at encoding textual information, traditional methods are still very competitive when it comes to encoding numeric and multimodal data. For our task, current LLMs perform worse than a more traditional XGBoost architecture that combines fundamental and macroeconomic data with high-density text-based embedding features. We investigate the degree to which the text encoding methodology affects performance and interpretability. The dataset reconstruction and model code from this paper is provided [^1] https://github.com/FelixDrinkall/credit-ratings-project.
Forecasting Credit Ratings: A Case Study where Traditional Methods Outperform Generative LLMs
Felix Drinkall*, Janet B. Pierrehumbert*‡, Stefan Zohren*† Department of Engineering Science, University of Oxford †The Alan Turing Institute ‡Faculty of Linguistics, University of Oxford felix.drinkall@eng.ox.ac.uk
1 Introduction
@@
Corporate credit ratings indicate a borrower’s ability to service its debt obligations and are a forward-looking measure of a company’s health Baresa et al. (2012). A company’s credit rating is significant since it affects the cost of raising capital, which in turn could finance future infrastructure to increase revenue or profitability. An optimistic rating can result in a virtuous cycle whereby it is easier to raise money and grow the business Cho et al. (2020), and a pessimistic rating can result in a vicious cycle in which competition can grow faster due to cheaper debt obligations. Knowing which cycle a company may enter can be advantageous to investors. Many major funds are also not allowed to own sub-prime assets, which makes forecasting a drop in credit rating very important so that the fund has more time to divest from the asset, which could result in a higher close price.
@@
Figure 1: Example of the best-performing feature - high-density clustering Drinkall et al. (2022). Each dot represents a sentence, and the colored areas representing high-density regions of the embedding space.
@@
Recently, there has been a surge of interest in text-based forecasting Xu and Cohen (2018); Yang et al. (2020); Nie et al. (2024). One reason for this trend is the progress that has been made in text modelling in general (Zoph et al., 2022; Touvron et al., 2023). Given that financial news is often first disseminated through written or spoken communications Boulland et al. (2016), rather than in numeric or tabular formats, there has been a hope that important information can be included in models sooner than was possible without using linguistic information. Another reason is that language can provide relevant context and forward-looking information, whereas financial numeric reporting alone is inherently retrospective. Contained within a company’s filings, text can provide insights about the future strategic direction of the company as well as historical information.
@@
The majority of text-based forecasting research has been focused on short text sequences, drawing primarily from sources such as social media Xu and Cohen (2018), news articles Zhang et al. (2018), and analyst recommendations Rekabsaz et al. (2017). In contrast, many fundamental financial documents, such as company filings, earnings call transcripts, and patents, are very long. Considering that the shorter texts often serve as summaries, reflections or commentaries on the detailed primary sources and that speed of information acquisition is essential in finance Rzayev et al. (2023), there must be more focus on text-based forecasting for longer text sequences. This paper evaluates the most effective ways to model longer text sequences within a text-based forecasting task.
@@
Linked to the recent progress of LLMs, there has been growing interest in applying these models in a variety of downstream applications Kaddour et al. (2023). It has been shown that as generative LLMs scale, they acquire abilities that were not present in smaller LLM variants Zoph et al. (2022), such as modular arithmetic Srivastava et al. (2023), NLU Hendrycks et al. (2021), commonsense reasoning Lin et al. (2022), fact-checking Rae et al. (2022) and so on. These abilities have been impressive, but it is best not to be over-optimistic. The lack of training data transparency associated with some of the best-performing LLMs means that we cannot be certain whether some of the performance gains are due to the memorisation of benchmarks being in the training datasets Bender et al. (2021); Sainz et al. (2023); Xu et al. (2024); Balloccu et al. (2024). Generative LLMs also seem to have a mediocre understanding of concepts like negation and complex logical reasoning Kassner and Schütze (2020); Lorge and Pierrehumbert (2023); Huang and Chang (2023); Truong et al. (2023). These limitations in the capabilities of LLMs could prove to be very consequential in financial contexts. In this paper, we test generative LLMs on a complex linguistic task, which has never been fully solved by human experts: credit rating forecasting. We show that while LLMs encode text-based information very well, they are not good at incorporating numeric information, and underperform a boosting-tree baseline.
The contributions of this paper are as follows:
@@
  • A benchmark of techniques for encoding long-sequence text in a forecasting task.
2 Related Work
2.1 Text-based forecasting
2.1.1 Encoding Text for Forecasting
@@
The predominant approach in text-based forecasting has focused on the extraction of interpretable features like sentiment and uncertainty scores Song and Shin (2019); An et al. (2023). Rule-based sentiment using diverse lexicons has dominated the literature Mohammad (2020); Kalamara et al. (2022); Barbaglia et al. (2023). Lexicons tailored to specific domains generally surpass broader lexicons in predictive tasks Loughran and McDonald (2011); Li et al. (2014). Nevertheless, lexicons overlook contextual nuance and inadequately address common linguistic phenomena like negation. To mitigate these limitations, efforts have been made to integrate more sophisticated sentiment classifiers An et al. (2023); Ayyappa et al. (2023). However, sentiment presupposes that important information can be encapsulated within a single dimension. To avoid an overly simple and prescriptive feature set, unsupervised methods have been used in feature exploration: TF-IDF Jones (1972), Latent Dirichlet Allocation (LDA) Wang et al. (2017); Kanungsukkasem and Leelanupab (2019). However, the arrival of contemporary topic models has gradually eclipsed LDA, fostering the adoption of transformer-derived topic models into forecasting tasks Drinkall et al. (2022).
@@
Recently, some studies have used the representations from encoder-based LLMs as features for text-based forecasting. LLMs exploit high-dimensional embeddings to capture the linguistic meaning of words Devlin et al. (2019); Radford et al. (2019) and sentences Reimers and Gurevych (2019), with representation dimensionality ranging from 384 Wang et al. (2020) to 5192 Touvron et al. (2023). Such dimensionality poses challenges when used with smaller datasets. Nonetheless, there has been some success in incorporating these methods into text-based forecasting Sawhney et al. (2020); Lee et al. (2023). However, the effectiveness of LLMs is often hampered by their limited context windows. Recent advancements have seen an increase in context window sizes, thanks in part to better GPU infrastructure making it computationally feasible, the implementation of attention sparsification techniques Tay et al. (2022), and positional encoding hacks Chen et al. (2023). There are also methods that combine the use of transformer-based LLMs with feature-based methods, such as topic clusters Grootendorst (2022); Drinkall et al. (2022), or emotions Liapis and Kotsiantis (2023).
2.1.2 Generative Multimodal Forecasting
@@
In addition to the adoption of encoder-based LLMs like BERT Devlin et al. (2019), generative LLMs have been used in text-based forecasting tasks. Generative LLMs use masked-self attention to model text in an autoregressive manner. The GPT Radford et al. (2019) and Llama Touvron et al. (2023) model families are part of the generative model class. LLMs have been used as a backbone model for generative time-series forecasting models Cao et al. (2023); Chang et al. (2023); Zhou et al. (2024); Liu et al. (2024), showing that an adapted generative language model can forecast weather, electricity and several other domains without relying on traditional text inputs. Liu et al. 2024 used eight text-based frames in order to create a general time-series modal that could be applied to several domains, and in so doing encoded both text and numerical information in a GPT-2-small model to generate the predicted time-series.
@@
Beyond the use of text-based frames in generative forecasting tasks, GPT4MTS Lee et al. (2023) encoded both news and time-series information before passing the concatenated input sequence through a pre-trained GPT-2 model. FinMA Xie et al. (2024) and PromptCast Xue and Salim (2024) evaluated the performance of LLMs on stock movement prediction by converting the time-series information into natural language and prompting the language model for the predicted direction. Yu et al. 2023 takes this further by passing exclusively text information into the prompt for a financial forecasting task. There has been little comparison between these generative methods and the more traditional discriminative methods when applied to multimodal information.
2.2 Credit Rating Prediction
@@
Research in Credit Rating Prediction (CRP) has tended to focus on predicting the absolute credit rating at time $t=0$ given the feature set $F_{t=0}$ Li et al. (2023); Galil et al. (2023); Tavakoli et al. (2023). This approach takes the perspective of the rating agencies and is useful for identifying anomalies where the existing credit rating classification appears to be implausible or inconsistent with current financial indicators Lokanan et al. (2019). However, predicting the absolute rating level is more simple and not as useful as predicting a future change. There is some limited research on Credit Rating Forecasting (CRF), where the target is the movement direction of the credit rating at time $t=1$. This task takes the perspective of the investor seeking to predict whether an asset is likely to be classified as more or less risky in the next time period, and is the task outlined in this paper.
@@
There have been some attempts to incorporate linguistic information into both corporate risk Fei et al. (2015); Cao et al. (2024) and default prediction Mai et al. (2019); Stevenson et al. (2021). Some papers have shown how text can help improve consumer credit lending Hurley and Adebayo (2016); Babaei and Giudici (2024). There has also been some attempts to include textual data in CRP and CRF tasks Chen and Chen (2022); Muñoz-Izquierdo et al. (2022); Tavakoli et al. (2023). The majority of the existing literature uses lexicons, keywords or sentiment to encode the text Kogan et al. (2009); Fei et al. (2015); Mai et al. (2019); Muñoz-Izquierdo et al. (2022); Chen and Chen (2022). There have been some studies that have utilized encoder-based LLM representations Stevenson et al. (2021); Tavakoli et al. (2023); Cao et al. (2024). There has been some work exploring how well generative models perform at assessing credit lending applications Babaei and Giudici (2024), and value at risk in general Cao et al. (2024), but there has been no work benchmarking how well modern generative LLMs perform on a CRF task. Understanding generative LLMs’ relative strengths relative to more traditional methods is an important contribution to the existing literature.
3 Dataset
@@
In part due to the lack of large open-source or readily available datasets with temporal metadata, most of the financial text-based forecasting studies have either focused on expensive proprietary datasets, or datasets spanning 2-3 years Xu and Cohen (2018); Soun et al. (2022), making results hard to replicate and potentially biased to a specific time. While temporal bias in language-based tasks is hard to avoid due to limited historical data Drinkall et al. (2024), we aim to reduce this by using a dataset spanning 23 years, increasing the models’ exposure to different economic contexts. The cost and lack of transparency of large datasets have hindered progress in the field and made it harder to build on promising work due to the difficulty of replicating results. As such, all data used in this paper is either open source or available with a WRDS subscription to enable effective dataset reconstruction. The data used in this paper is from US-based companies.
3.0.1 Credit ratings (C)
@@
For the credit ratings, we used the Compustat Capital IQ dataset [^2] Credit Ratings: https://tinyurl.com/r4urtkc5, using Standard & Poors’ (S&P) ratings. These ratings cover the period from 1978 to 2017. S&P routinely assesses and assigns credit ratings to companies. Our paper predicts changes to the long-term credit ratings. Notably, we incorporate historical ratings from preceding quarters into our prediction models, acknowledging the distinct implications of a top-rated company (AAA) being downgraded compared to a lower-rated one (CC) experiencing a similar decline.
3.0.2 SEC filings
@@
This paper uses 10-Q and 10-K filings available in the SEC’s EDGAR database [^3] Filings Database: https://tinyurl.com/3rdn7hrx to provide both textual context. They were chosen for their consistent structure which aids homogenous feature extraction. While most of the content in these filings is comprised of indexing, tables, and introductory text, we’re interested in the parts that offer insights into a company’s future financial health. As such, we’ve focused on the Management’s Discussion and Analysis of Financial Condition and Results of Operations (MDA) section. We extract the MDA sections from all SEC filings - using the SEC-API [^4] Filings API: https://sec-api.io/ - for which we had credit rating data, spanning from Q1-1994 to Q2-2017. The API returns cleaned text, but we clean the text further by removing the remaining HTML, links and excessive spaces.
3.0.3 Fundamental data (F)
@@
S&P emphasizes two components in their credit rating methodology: the financial and business risk profiles Gillmor (2015). While the text from the MDA section provides some insight into the qualitative business risk profile, numerical fundamental data is important to assess the financial health of a company. For the fundamental data, we use the Compustat Quarterly Fundamentals dataset [^5] Fundamental Data:https://tinyurl.com/4ca8ddst. The variables selected are outlined in Appendix C. These variables were consistently reported for all the companies under consideration. Ideally, we would incorporate a broader range of fundamental variables, but expanding the variable set would result in fewer samples with complete data, thus limiting the scope of our analysis.
3.0.4 Macroeconomic data (M)